Please help with my question, I can catch on fast if I am explained how to do it
Find the equation of the line, both in point-slope and slope intercept form going through the point (3,5) and parallet to the line with the equation
y=-x+5

Answer by ankor@dixie-net.com(22740)

Find the equation of the line, both in point-slope and slope intercept
form going through the point (3,5) and parallel to the line with the
equation: y=-x+5
This is in the slope/intercept form
Parallel lines have the same slope, therefore the parallel line
y = -1x + b
Given points x=3, y=5, find b
5 = -1(3) + b
5 + 3 = b
b = 8
y = -x + 8 is the parallel line going thru 3,5
:
The point-slope form y - y1 = m(x - x1)
therefore
y - 5 = -1(x - 3)